A system is given by:      

yt=5xt+1-3

Is the system BIBO stable? Justify

Is the system memoryless? Explain

Is the system causal? Explain

Is the system time invariant? Justify

Is the system linear? Justify

What is the impulse response h(t) of the system?

Is the system internally stable? If you could not figure out h(t) from part f, use the h(t) from the problem below.

ht=5e-tut-u(t-4)
Answer #1

bol: → Jet) 52(t+1-3 (a) BIBO Stability: – det act) = 4(6) bolended input ylt) 5 () 33:46 H€) (54C+0-3] act) YCt? 1 -3 1 FoomFes linearity : Let x 20,(+) => ax(+) 41(t) = 5a 2(t+1)-3 32(t) = b 20 €) 21t) — 56X6H1 -3 Let 33 () = alt) + 620(+) (t) + 53